What are the dimensions of a rectangle that has a perimeteer of 42 meters and an area of 90 square meters.

HELP!

Answer by checkley75(3666)

2L+2W=42
2L=42-2W
L=42/2-2W/2
L=21-W
LW=90
(21-W)W=90
21W-W^2=90
-W^2+21W-90=0
W^2-21W+90=0
(W-15)(W-6)=0
W-15=0
W=15 ANSWER.
W-6=0
W=6 ANSWER.
PROOF
2*6=2*15=42
12+30=42
42=42